# Instructions for Authors: Title in Title Case

### v1

@@ B.4.5 Results of Extra Ablation Study Experiments on ML-100K Dataset
Table 17: Comparison made between original algorithms and baselines using average method when training on ML-100K dataset. Subtable 17(a) is the performance made by original ones, which use whole data to train. Subtable 17(b) is the performance made by algorithms using average method (only baselines, excluding SBUCB and IBCB), which use every episode data training distinct models, and final model uses simple average method to average these models’ parameters.
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.891±0.006 0.193±0.005 /
B-IRL 0.830±0.010 0.189±0.005 43.669±1.282
BaseSVC / 0.189±0.005 162.767±12.135
B-ICB-S 0.746±0.032 0.187±0.007 147.474±4.633
IBCB (Ours) 0.870±0.014 0.192±0.006 15.023±3.854
(a) results of the original experiment
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.891±0.006 0.193±0.005 /
B-IRL 0.835±0.007 0.190±0.005 85.352±9.192
BaseSVC / 0.189±0.004 20.619±0.377
B-ICB-S 0.336±0.076 0.129±0.019 214.288±2.558
IBCB (Ours) 0.870±0.014 0.192±0.006 19.514±4.891
(b) results of averaging $N$ baseline model’s parameters
@@
Table 18: Comparison made between larger batch size $B$ ($B=2000$) and smaller batch size $B$ ($B=500$) when training & testing on ML-100K dataset. Subtable 18(a) is the experiment results obtained when $B=2000$. Subtable 18(b) is the experiment results obtained when $B=500$.
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.896±0.001 0.194±0.005 /
B-IRL 0.816±0.023 0.192±0.005 52.485±1.241
BaseSVC / 0.191±0.004 230.26±12.606
B-ICB-S 0.685±0.028 0.188±0.006 144.697±7.972
IBCB (Ours) 0.851±0.016 0.196±0.007 15.083±2.036
(a) results of setting batch size $B=2000$
Algorithm OL-Fitness BT-AR Train Time (sec.)
SBUCB 0.89±0.012 0.193±0.006 /
B-IRL 0.845±0.007 0.190±0.006 65.646±1.378
BaseSVC / 0.189±0.005 266.538±12.153
B-ICB-S 0.716±0.039 0.187±0.007 138.616±8.982
IBCB (Ours) 0.862±0.019 0.191±0.005 26.255±10.370
(b) results of setting batch size $B=500$
@@
Firstly, since we assume that training period can be described as a period containing $N$ episodes’ training data, we want to know how this $N$’s setting influence the performance of baselines. Specifically, we divided the total $N$ episodes to individually train corresponding IRL or BC models. Finally, we averaged the parameters of the $N$ models obtained from each episode’s training to form the final test model. In this setup, the training time equates to the cumulative time for training $N$ models (disregarding multi-threading, assuming that each time we collect data from an episode, we train a model sequentially). The experiment was conducted using the ML-100K dataset, as we aimed to investigate the practical feasibility. Detailed experimental results can be found in Table 17. From the experimental results, it is evident that training IRL or BC models separately for each episode does yield certain differences in comparison to the unified training approach. The B-IRL model demonstrates a slight improvement in performance (e.g. OL-Fitness), albeit with a significant increase in training time. This could be attributed to the necessity of restarting training each time due to the inability to iteratively utilize previously acquired parameters when new data is collected. Of particular interest is the significant decline in performance for B-ICB. This discrepancy is likely linked to the inconsistency in retaining the original expert model’s (i.e. SBUCB) parameters within each episode, which contradicts B-ICB’s assumption of evolving data within every single step. Additionally, due to the necessity for repeated restarts during training, the overall training time becomes extended. As for BaseSVC, the reduced training sample size in each episode leads to notably shorter training time for individual runs (attributed to its near-quadratic time complexity). While this reduces overall training time, the efficacy is comparatively diminished compared to the direct training using the entire dataset. Ultimately, even in comparison with the models subjected to parameter averaging, the performance of IBCB remains superior overall. This underscores the effectiveness of using episodes as evolutionary segments in IBCB’s training process in real-world datasets.
@@
Secondly, we wanted to work out the influence of batch size $B$’s setting. Throughout this ablation study experiment, we controlled the total number of training and testing samples and fixed the context, ensuring that the product of the total number of episodes $N$ and the batch size $B$ within each episode remains constant. Based on this foundation, we adjusted the batch size $B$. Specifically, the original values were $N=20$ and $B=1000$. Extra experiments included scenarios with larger $B$, denoted as $N=10$ and $B=2000$ , with corresponding results displayed in Table 18(a). Likewise, there were scenarios with smaller $B$, represented as $N=40$ and $B=500$, with results exhibited in Table 18(b). From the results of these additional experiments and comparisons with the original experiments, we deduced that, while controlling the total number of training and testing samples and context, a larger $B$ yields shorter overall training time and better average reward performance on the test set. However, the fitting to the environment is reduced due to the decreased value of $N$, implying that more data for training comprises novice expert’s evolutionary history, hence fostering more exploration. With reduced $N$ and increased $B$, the frequency of evolution decreases, making adaptation to the original training data easier, thereby shortening training time. In such cases, the estimations of various baselines and IBCB regarding the environment are not as precise as those under smaller $B$. Appropriately crafting exploration items can result in higher rewards. Conversely, a smaller $B$ leads to a better fit with the environment, albeit resulting in a longer training time and inferior average reward performance on the test set, which aligns with the experimental results.

### v2

@@ 5.6.7 Extra Ablation Study Experiments of Batch Test period on ML-100K Dataset
TABLE XIX: Comparison made between original algorithms and baselines using average method when training on ML-100K dataset. Upper Subtable is the performance made by original ones, which use whole data to train. Lower Subtable is the performance made by algorithms using average method (only baselines, excluding SBUCB and IBCB), which use every episode data training distinct models, and final model uses simple average method to average these models’ parameters.
Algorithm OL-Fitness BT-Fitness BT-AR Train Time (sec.)
SBUCB 0.891±0.006 1 0.193±0.005 /
B-IRL 0.830±0.010 0.864±0.008 0.189±0.005 43.669±1.282
BaseSVC / 0.804±0.012 0.189±0.005 162.767±12.135
B-ICB-S 0.746±0.032 0.767±0.028 0.187±0.007 147.474±4.633
IBCB (Ours) 0.862±0.023 0.925±0.019 0.192±0.005 4.036±0.111
Algorithm OL-Fitness BT-Fitness BT-AR Train Time (sec.)
SBUCB 0.891±0.006 1 0.193±0.005 /
B-IRL 0.835±0.007 0.868±0.008 0.190±0.005 85.352±9.192
BaseSVC / 0.748±0.009 0.189±0.004 20.619±0.377
B-ICB-S 0.336±0.076 0.262±0.064 0.129±0.019 214.288±2.558
IBCB (Ours) 0.862±0.023 0.925±0.019 0.192±0.005 4.036±0.111
@@
TABLE XX: Comparison made between larger batch size $B$ ($B=2000$) and smaller batch size $B$ ($B=500$) when training & testing on ML-100K dataset. Upper Subtable is the experiment results obtained when $B=2000$. Lower Subtable is the experiment results obtained when $B=500$.
Algorithm OL-Fitness BT-Fitness BT-AR Train Time (sec.)
SBUCB 0.896±0.001 1 0.194±0.005 /
B-IRL 0.816±0.023 0.866±0.014 0.192±0.005 52.485±1.241
BaseSVC / 0.807±0.014 0.191±0.004 230.26±12.606
B-ICB-S 0.685±0.028 0.707±0.052 0.188±0.006 144.697±7.972
IBCB (Ours) 0.861±0.010 0.886±0.014 0.192±0.004 2.907±0.261
Algorithm OL-Fitness BT-Fitness BT-AR Train Time (sec.)
SBUCB 0.890±0.012 1 0.193±0.006 /
B-IRL 0.845±0.007 0.877±0.009 0.190±0.006 65.646±1.378
BaseSVC / 0.815±0.006 0.189±0.005 266.538±12.153
B-ICB-S 0.716±0.039 0.723±0.049 0.187±0.007 138.616±8.982
IBCB (Ours) 0.835±0.023 0.833±0.019 0.194±0.006 4.296±0.128
@@
Firstly, since we assume that training period can be described as a period containing $N$ episodes’ training data, we want to know how this $N$’s setting influence the performance of baselines. Specifically, we divided the total $N$ episodes to individually train corresponding IRL or BC models. Finally, we averaged the parameters of the $N$ models obtained from each episode’s training to form the final test model. In this setup, the training time equates to the cumulative time for training $N$ models (disregarding multi-threading, assuming that each time we collect data from an episode, we train a model sequentially). The experiment was conducted using the ML-100K dataset, as we aimed to investigate the practical feasibility. Detailed experimental results can be found in Table XIX. From the experimental results, it is evident that training IRL or BC models separately for each episode does yield certain differences in comparison to the unified training approach. The B-IRL model demonstrates a slight improvement in performance (e.g. OL-Fitness), albeit with a significant increase in training time. This could be attributed to the necessity of restarting training each time due to the inability to iteratively utilize previously acquired parameters when new data is collected. Of particular interest is the significant decline in performance for B-ICB. This discrepancy is likely linked to the inconsistency in retaining the original expert model’s (i.e. SBUCB) parameters within each episode, which contradicts B-ICB’s assumption of evolving data within every single step. Additionally, due to the necessity for repeated restarts during training, the overall training time becomes extended. As for BaseSVC, the reduced training sample size in each episode leads to notably shorter training time for individual runs (attributed to its near-quadratic time complexity). While this reduces overall training time, the efficacy is comparatively diminished compared to the direct training using the entire dataset. Ultimately, even in comparison with the models subjected to parameter averaging, the performance of IBCB remains superior overall. This underscores the effectiveness of using episodes as evolutionary segments in IBCB’s training process in real-world datasets.
@@
Secondly, we wanted to work out the influence of batch size $B$’s setting. Throughout this ablation study experiment, we controlled the total number of training and testing samples and fixed the context, ensuring that the product of the total number of episodes $N$ and the batch size $B$ within each episode remains constant. Based on this foundation, we adjusted the batch size $B$. Specifically, the original values were $N=20$ and $B=1000$. Extra experiments included scenarios with larger $B$, denoted as $N=10$ and $B=2000$ , with corresponding results displayed in upper subtable of Table XX. Likewise, there were scenarios with smaller $B$, represented as $N=40$ and $B=500$, with results exhibited in lower subtable of Table XX. From the results of these additional experiments and comparisons with the original experiments, we deduced that, while controlling the total number of training and testing samples and context, a larger $B$ yields shorter overall training time and better fitness on both environment and expert algorithm’s parameters. Therefore, the fitting to the environment and expert is reduced due to the increased value of $N$, implying that more data for training comprises novice expert’s evolutionary history, hence fostering more exploration. With reduced $N$ and increased $B$, the frequency of evolution decreases, making adaptation to the original training data easier, thereby shortening training time. We found that the smaller $B$, the larger the total number of episodes $N$, but the extent of subsequent episode exploration gradually becomes smaller, so IBCB may tend to overfit. On the contrary, when the $B$ is larger, the smaller the total number of episodes $N$, so the environment will encourage more exploration, and IBCB can fit better environment and expert algorithm’s parameters. In such cases, IBCB shows preciser performance against other baselines under larger $B$. Appropriately crafting exploration items can result in higher rewards.
6 Conclusion
This paper aims to address the problem of efficiently learning behaviours from the expert’s evolution history. Specifically, we propose an inverse batched contextual bandit model called IBCB. The proposed IBCB gives a unified framework for both deterministic and randomized bandit policies, and solves the problem of learning from evolution history with inaccessible rewards through a simple quadratic programming problem. Experimental results demonstrate the effectiveness and efficient training of IBCB on both synthetic and real-world scenarios. Besides, IBCB can also be generalized to out-of-distribution and contradictory data scenarios with great robustness. When it comes to limitations, IBCB needs to be further explored in experiments related to fairness, such as more random action selection methods, fairness assumptions based on specific distributions and other related settings. In addition, this article mainly discusses the UCB-like gambling machine framework. In the future work, we will focus on better random optimization of IBCB.